# Rodrigues Network for Learning Robot Actions

### v1

@@ 4.2 Joint Layer
While the Rodrigues Layer updates the link features, we still need a learnable mechanism to update the joint features. The Joint Layer computes the output joint features $\{\Theta_{j}^{\text{out}}\}$ from the input joint features $\{\Theta_{j}^{\text{in}}\}$ and link features $\{F_{l}^{\text{in}}\}$. For each joint $J_{j}$, we retrieve the feature of its child link $F^{\text{in}}_{\text{c}_{j}}$, apply a joint-specific linear transformation, and add it to the joint’s existing feature:
@@
The transformations $\text{Linear}_{j}:\mathbb{R}^{C_{L}\times 4\times 4}\to\mathbb{R}^{C_{J}}$ are independently learned for each joint, allowing the model to capture joint-specific information. This operator enables the propagation of information from links to joints while preserving the locality of the robot’s articulation structure.
@@ 4.3 Other components and overall architecture
While the Rodrigues Layer and Joint Layer leverage the spatial locality inherent in articulated structures, they restrict information flow to consecutive links and joints. To enable direct communication across all links, we incorporate a Self-attention Layer. In this layer, each link feature is first projected into a token using a linear transformation. These tokens then interact through multi-head self-attention, allowing the model to aggregate information from all links regardless of their spatial distance. The attended tokens are subsequently projected back to the link feature space, followed by residual addition and layer normalization.
@@
To capture and utilize global context, we optionally introduce a global token $G$. This token represents a learnable global feature and participates in self-attention alongside the link tokens. It is processed through its own projection, residual, and normalization steps, and joins the link tokens during multi-head self-attention. The global token enables the network to store and propagate task-wide information that is not tied to any specific joint or link. We optionally enable the global token in tasks that require predicting global outputs, such as base link pose estimation.
@@
We combine the above three components into a unified module called the Rodrigues Block. Each Rodrigues Block takes as input the link features, joint features, and an optional global token. It sequentially applies the Rodrigues Layer, Joint Layer, and Self-attention Layer to produce updated link features, joint features, and a global token. By stacking multiple such blocks, we construct the full Rodrigues Network (RodriNet), enabling deep, hierarchical reasoning over articulated structures. The overall architecture is drawn in Figure 2.
@@ 5 Experiments
We evaluate our Rodrigues Network on a set of different tasks, ranging from forward kinematics and motion prediction (Section 5.1), to imaging learning in robotics (Section 5.2), to hand pose estimation for vision and graphics (Section 5.3). The experiments focus on two main questions on action learning: (i) Does the structural prior of articulated kinematics make the Rodrigues Network better at representing robot actions? (ii) Can this inductive bias improve the understanding and prediction of robot actions in realistic task scenarios?
@@ 5.1 Toy experiments on kinematics and motion
Figure 4: Visualization of forward kinematics prediction on an example configuration. Errors are plotted on each link with color scales, with darker colors indicating larger errors. MLP learns better positions but much worse orientations than GCN, leading to more aligned links but darker colors.
@@
As discussed in Section 3.1, forward kinematics maps the configuration of an articulated robot (including the root pose and joint angles) to the pose of each link. To effectively control a robot, a neural network should, at a minimum, be capable of learning this mapping. To evaluate whether the Neural Rodrigues Operator possesses strong representational capacity for kinematic modeling, we construct a specialized Rodrigues Network consisting solely of Rodrigues Layers and compare it to other neural backbones in their ability to fit the forward kinematics of a single robot. We conduct this experiment on the LEAP Hand [42], a free-floating dexterous robotic hand with 16 joints and 17 links. The network takes as input the configuration $(T,R,\theta)$ and predicts the pose matrices for all 17 links. For model hyperparameters and training settings, please refer to the supplementary material.
@@
Table 1: Motion prediction in Cartesian space with trainset size $=10^{5}$.
Backbone $\text{Error}_{T}$ (mm) $\text{Error}_{R}$ () $\text{Error}_{\theta}$ () MSE ($1\mathrm{e}{-6}$) Train MSE ($1\mathrm{e}{-6}$)
MLP 3.49 $\pm$0.05 0.46 $\pm$0.05 0.17 $\pm$0.00 22.52 $\pm$0.95 12.47 $\pm$0.73
GCN 3.55 $\pm$0.05 0.48 $\pm$0.05 0.17 $\pm$0.01 18.52 $\pm$1.74 13.68 $\pm$1.87
BoT 2.92 $\pm$0.04 0.46 $\pm$0.04 0.15 $\pm$0.01 15.72 $\pm$1.21 13.04 $\pm$1.41
Transformer 2.89 $\pm$0.06 0.41 $\pm$0.06 0.14 $\pm$0.01 12.86 $\pm$1.25 10.50 $\pm$1.21
Rodrigues 1.21 $\pm$0.04 0.16 $\pm$0.04 0.06 $\pm$0.00 2.56 $\pm$0.39 1.93 $\pm$0.34
@@
Although fitting forward kinematics highlights the benefits of incorporating kinematics-inspired inductive bias, the task itself has limited practical value. In real-world robot learning scenarios, neural backbones typically process observations in 3D Cartesian space (e.g., point clouds) and output control commands as target joint angles. To evaluate a network’s ability to bridge these modalities, we propose a Cartesian motion prediction task that challenges the model to reason about structured motion in Cartesian space while expressing its predictions in joint angle space. This task uses a 6-DoF UR5 robotic arm. We begin by randomly sampling two end-effector poses in 3D space and interpolating between them (linearly for translation and spherically for rotation) to generate a 16-frame Cartesian trajectory. Using inverse kinematics, we convert this trajectory into corresponding joint configurations. The network is given the first 8 frames of joint angles and tasked with predicting the remaining 8, also in joint space. For details on training and model parameter count control, please refer to the supplementary.
As shown in Table 1, the Rodrigues Network achieves the lowest training loss and test errors when trained on a pre-collected dataset of $10^{5}$ trajectories. Notably, its test MSE is lower than the train MSE of all baseline models, indicating that the Rodrigues Network not only fits the data more effectively but also generalizes better without overfitting. Figure 5(b) further shows that our model consistently outperforms baselines across different training set sizes. In Figure 5(a), we visualize the end-effector trajectories predicted by each network for a single example; our model’s trajectory aligns most closely with the ground truth. These results confirm that the Rodrigues Network is more effective at bridging Cartesian space and joint angle space than conventional architectures.
@@ 5.2 Robotic manipulation with imitation learning
Table 2: Baseline comparisons on the imitation learning benchmark.
@@
Next, we evaluate whether our method benefits realistic robotic applications. We integrate the Rodrigues Network as a backbone into the Diffusion Policy [14], one of the state-of-the-art imitation learning frameworks, and test on a manipulation benchmark in simulation.
Benchmark We construct a suite of five manipulation tasks from ManiSkill [35] using a 7-DoF Franka arm with a 1-DoF Panda gripper, simulated in SAPIEN [52]. For each task, we collect 100–500 demonstration trajectories using motion planning, with each trajectory spanning 200 steps. During training, the network receives proprioception and object state as input and outputs relative joint offsets used for PD control. Performance is measured by running 100 evaluation rollouts in simulation, and all models are trained with 5 random seeds to report the mean and standard deviation of success rates.
Methods We adopt Diffusion Policy [14] as our learning framework, using a 2-frame observation history and predicting 16 future steps, of which 8 are executed during deployment. To isolate the impact of the neural backbone, we keep the outer framework and all other components fixed, modifying only the denoising network. Baseline architectures include the U-Net and Transformer designs from the original Diffusion Policy paper. Our method replaces these with the Rodrigues Network, which takes the current observation, denoising timestep, and a noisy action as inputs and predicts the corresponding action noise. The gripper output is handled via the global token. All models are tuned to have approximately 17 million parameters for a fair comparison. Implementation details are provided in the supplementary material.
Results and analysis As shown in Table 2, Diffusion Policy [14] with the Rodrigues Network backbone achieves overall state-of-the-art performance, demonstrating that our kinematics-inspired inductive bias improves imitation learning in realistic robotic tasks. In particular, we observe substantial gains in PickCube and StackCube, and comparable or slightly better performance in PushCube, PegInsertionSide, and PlugCharger. These results suggest that the benefits of the Rodrigues Decoder are task-dependent. PushCube is relatively simple, leading all backbones to achieve near-perfect success. In contrast, PegInsertionSide and PlugCharger involve complex contact dynamics, such as inserting a peg into a hole, that likely require tactile or force feedback, making the neural backbone less of a limiting factor. In summary, enhancing the backbone architecture with our kinematics-inspired inductive bias can yield significant gains when the network is the performance bottleneck.
@@
Table 3: Baseline comparisons on the FreiHAND dataset [59]. We use the standard protocol and report metrics on 3D joint and 3D mesh accuracy. PA-MPVPE and PA-MPJPE numbers are in mm.
@@
I2L-MeshNet [32] 7.4 7.6 0.681 0.973
Pose2Mesh [15] 7.7 7.8 0.674 0.969
I2UV-HandNet [10] 6.7 6.9 0.707 0.977
METRO [28] 6.5 6.3 0.731 0.984
Tang et al. [47] 6.7 6.7 0.724 0.981
Mesh Graphormer [29] 5.9 6.0 0.764 0.986
MobRecon [11] 5.7 5.8 0.784 0.986
AMVUR [24] 6.2 6.1 0.767 0.987
HaMeR [37] 6.0 5.7 0.785 0.990
HaMeR (Reproduced) [37] 6.2 5.9 0.774 0.989
@@
Finally, we demonstrate that our Rodrigues Network is not limited to robot learning but also applies to the animatable models in computer vision and graphics, where articulated actors commonly exist, such as human-like characters and hands. Specifically, we apply our method to 3D hand reconstruction from single-view RGB images, which involves predicting the rotations and positions of hand joints based on the kinematic structure defined by the MANO [40] hand kinematics.
Our network builds upon HaMeR [37] by replacing its vanilla transformer with the proposed Rodrigues Network (with modifications to suit MANO’s configuration representation). Additionally, we introduce a cross-attention layer following the self-attention layer to enable interactions between joint and link features and the input image tokens. Our network achieves a notable performance improvement while significantly reducing the number of trainable parameters (39.5M vs. ours: 10.7M). Further architectural details are provided in the supplementary material.
Table 3 presents the quantitative results of our method. We follow the evaluation protocol and metrics established by HaMeR [37], and report performance on the FreiHand [59] dataset. For reference, we also include the results of our own reproduction of the HaMeR model.
@@ 6 Conclusions and discussions
First, the Rodrigues Network focuses on leveraging the kinematic structure at the level of joints, while overlooking the geometry of the links, such as their lengths and shapes. More effectively incorporating link information and geometry awareness could provide a more comprehensive representation of the kinematic structure, potentially improving performance on tasks that require precise contact reasoning. Second, the current Neural Rodrigues Operator is designed for rotational joints; extending it to handle both rotational and translational joints could broaden the network’s applicability to a wider range of robotic systems. Third, our robot learning experiments focused on imitation learning. An important direction for future work is to investigate how the network performs under reinforcement learning (RL), which could enhance its generality and effectiveness in closed-loop learning scenarios.
@@ 7 Acknowledgements
Leonidas Guibas, Yang You and Congyue Deng acknowledge support from the Toyota Research Institute University 2.0 Program, ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a gift from the Flexiv corporation. Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University. Pieter Abbeel holds concurrent appointments as a professor at UC Berkeley and as an Amazon Scholar, and this paper is not associated with Amazon.
@@
This supplementary material provides additional details to support the experiments presented in the main paper. Section A describes how we adapt the Neural Rodrigues Operator for animated characters, as used in the 3D hand reconstruction experiment. Section B outlines our experimental settings, including data collection, training procedures, evaluation protocols, and all training hyperparameters. Section C details the architecture and configurations of both baseline models and our proposed method across all experiments. These details are provided to facilitate reproducibility of our results. We will release our code, datasets, and pretrained checkpoints upon acceptance to further support the community in reproducing and building upon our work.
Additionally, Section D presents further experiments designed to investigate the properties of our model and provide additional support for the claims made in the main paper. Finally, Section E describes our CUDA implementation of the Multi-Channel Rodrigues Operator, which accelerates training and inference.

### v2

@@ 4.2 Joint Layer
While the Rodrigues Layer updates the link features, we still need a learnable mechanism to update the joint features. The Joint Layer computes its output joint features $\{\Theta_{j}^{\text{out}}\}$ from its own input joint features $\{\Theta_{j}^{\text{in}}\}$ and link features $\{F_{l}^{\text{in}}\}$. For each joint $J_{j}$, we retrieve the feature of its child link $F^{\text{in}}_{\text{c}_{j}}$, apply a joint-specific linear transformation, and add it to the joint’s existing feature:
@@
The transformations $\text{Linear}_{j}:\mathbb{R}^{C_{L}\times 4\times 4}\to\mathbb{R}^{C_{J}}$ are independently learned for each joint, allowing the model to capture joint-specific information. This layer updates the joint features and leaves the link features unchanged.
@@ 4.3 Other components and overall architecture
While the Rodrigues Layer and Joint Layer leverage the spatial locality inherent in articulated structures, they restrict information flow to consecutive links and joints. To enable direct communication across all links, we incorporate a Self-attention Layer. In this layer, each link feature is first projected into a token using a linear transformation. These tokens then interact through multi-head self-attention, allowing the model to aggregate information from all links regardless of their spatial distance. The attended tokens are subsequently projected back to the link feature space, followed by residual addition and layer normalization. Joint features are left unchanged.
@@
To capture and utilize global context, we optionally introduce a global token $G$. This token represents a learnable global feature and only participates in self-attention alongside the link tokens. It is processed through its own projection, residual, and normalization steps, and joins the link tokens during multi-head self-attention. The global token enables the network to store and propagate task-wide information that is not tied to any specific joint or link. We optionally enable the global token in tasks that require predicting global outputs, such as base link pose estimation.
@@
We combine the above three components into a unified module called the Rodrigues Block. Each Rodrigues Block takes as input the link features, joint features, and an optional global token. It sequentially applies the Rodrigues Layer, Joint Layer, and Self-attention Layer to produce updated link features, joint features, and a global token, where each layer’s outputs serves as the next layer’s inputs. By stacking multiple such blocks sequentially, we construct the full Rodrigues Network (RodriNet), enabling deep, hierarchical reasoning over articulated structures. The overall architecture is drawn in Figure 2. Refer to Section B of the supplementary for details on computing the first-layer features and task-specific outputs.
@@ 5 Experiments
We evaluate our Rodrigues Network on a set of different tasks, ranging from forward kinematics and motion prediction (Section 5.1), to imitation learning in robotics (Section 5.2), to hand pose estimation for vision and graphics (Section 5.3). The experiments focus on two main questions on action learning: (i) Does the structural prior of articulated kinematics make the Rodrigues Network better at representing robot actions? (ii) Can this inductive bias improve the understanding and prediction of robot actions in realistic task scenarios? Additional studies on ablations and hyperparameter sensitivity are provided in the supplementary material.
@@ 5.1 Toy experiments on kinematics and motion
Figure 4: Visualization of forward kinematics prediction on an example configuration. Errors are plotted on each link with color scales, with darker colors indicating larger errors.
@@
As discussed in Section 3.1, forward kinematics maps the configuration of an articulated robot (including the root pose and joint angles) to the pose of each link. To effectively control a robot, a neural network should, at a minimum, be capable of learning this mapping. To evaluate whether the Neural Rodrigues Operator possesses strong representational capacity for kinematic modeling, we construct a specialized Rodrigues Network consisting solely of Rodrigues Layers and compare it to other neural backbones in their ability to fit the forward kinematics of a single robot. We conduct this experiment on the LEAP Hand (Shaw et al., 2023), a free-floating dexterous robotic hand with 16 joints and 17 links. The network takes as input the configuration $(T,R,\theta)$ and predicts the pose matrices for all 17 links. For model hyperparameters and training settings, please refer to Sections C.1 and D.1 of the supplementary material.
@@
Table 1: Motion prediction in Cartesian space with trainset size $10^{5}$.
Backbone $\text{Error}_{T}$ ($m$) $\text{Error}_{R}$ () $\text{Error}_{\theta}$ () MSE ($1\mathrm{e}{-6}$) Train MSE ($1\mathrm{e}{-6}$)
MLP 3.49 $\pm$0.33 0.46 $\pm$0.05 0.17 $\pm$0.00 22.52 $\pm$0.95 12.47 $\pm$0.73
GCN 3.55 $\pm$0.44 0.48 $\pm$0.05 0.17 $\pm$0.01 18.52 $\pm$1.74 13.68 $\pm$1.87
BoT 2.92 $\pm$0.29 0.46 $\pm$0.04 0.15 $\pm$0.01 15.72 $\pm$1.21 13.04 $\pm$1.41
Transformer 2.89 $\pm$0.45 0.41 $\pm$0.06 0.14 $\pm$0.01 12.86 $\pm$1.25 10.50 $\pm$1.21
Rodrigues 1.21 $\pm$0.17 0.16 $\pm$0.04 0.06 $\pm$0.00 2.56 $\pm$0.39 1.93 $\pm$0.34
@@
Although fitting forward kinematics highlights the our operator’s representation capacity for kinematics, the task itself has limited practical value. In real-world robot learning scenarios, neural backbones typically process observations in 3D Cartesian space (e.g., point clouds) and output control commands as target joint angles. To evaluate a network’s ability to bridge these modalities, we propose a Cartesian motion prediction task that challenges the model to reason about structured motion in Cartesian space while expressing its predictions in joint angle space. This task uses a 6-DoF UR5 robotic arm. We begin by randomly sampling two end-effector poses in 3D space and interpolating between them (linearly for translation and spherically for rotation) to generate a 16-frame Cartesian trajectory. Then, we use inverse kinematics to convert this trajectory into corresponding joint configurations. The network is given the first 8 frames of joint angles and tasked with predicting the remaining 8, also in joint space. Refer to Sections C.2 and D.2 of the supplementary for details on training, model parameter count control, and runtime comparisons.
As shown in Table 1, the Rodrigues Network achieves the lowest training loss and test errors when trained on a pre-collected dataset of $10^{5}$ trajectories. Notably, its test MSE is lower than the train MSE of all baseline models, indicating that the Rodrigues Network not only fits the data more effectively but also generalizes better without overfitting. Figure 5(b) further shows that our model consistently outperforms baselines across different training set sizes. In Figure 5(a), we visualize the end-effector trajectories predicted by each network for a single example; our model’s trajectory aligns most closely with the ground truth. These results confirm that the Rodrigues Network is more effective at bridging Cartesian space and joint angle space than conventional architectures. Please refer to Section C.2 of the supplementary for metric definitions.
@@ 5.2 Robotic manipulation with imitation learning
Table 2: Baseline comparisons on the imitation learning benchmark. Simulated success rate.
@@
Next, we evaluate whether our method benefits realistic robotic applications. We integrate the Rodrigues Network as a backbone into the Diffusion Policy (Chi et al., 2023), one of the state-of-the-art imitation learning frameworks, and test on a manipulation benchmark in simulation.
Benchmark We construct a suite of five manipulation tasks from ManiSkill (Mu et al., 2021) using a 7-DoF Franka arm with a 1-DoF Panda gripper, simulated in SAPIEN (Xiang et al., 2020). For each task, we collect 100–500 demonstration trajectories using motion planning, with each trajectory spanning 200 steps. During training, the network receives proprioception and object state as input and outputs relative joint offsets used for PD control. Performance is measured by running 100 evaluation rollouts in simulation, and all models are trained with 5 random seeds to report the mean and standard deviation of success rates.
Methods We adopt Diffusion Policy (Chi et al., 2023) as our learning framework, using a 2-frame observation history and predicting 16 future steps, of which 8 are executed during deployment. To isolate the impact of the neural backbone, we keep the outer framework and all other components fixed, modifying only the denoising network. Baseline architectures include the U-Net and Transformer designs from the original Diffusion Policy paper. Our method replaces these with the Rodrigues Network, which takes the current observation, denoising timestep, and a noisy action as inputs and predicts the corresponding action noise. The gripper output is handled via the global token. All models are tuned to have approximately 17 million parameters for a fair comparison. Implementation details are provided in Section D.3 of the supplementary material.
Results and analysis As shown in Table 2, Diffusion Policy (Chi et al., 2023) with the Rodrigues Network backbone achieves overall state-of-the-art performance, demonstrating that our kinematics-inspired inductive bias improves imitation learning in realistic robotic tasks. In particular, we observe substantial gains in PickCube and StackCube, and comparable or slightly better performance in PushCube, PegInsertionSide, and PlugCharger. These results suggest that the benefits of the Rodrigues Decoder are task-dependent. PushCube is relatively simple, leading all backbones to achieve near-perfect success. In contrast, PegInsertionSide and PlugCharger involve complex contact dynamics, such as inserting a peg into a hole, where tactile or force feedback would be beneficial. However, these sensor inputs are not provided in the environment, making the neural backbone less of a limiting factor. In summary, enhancing the backbone architecture with our neural operator can yield significant gains when the network is the performance bottleneck.
@@
Table 3: Baseline comparisons on the FreiHAND dataset. We use the standard protocol and report metrics on 3D joint and 3D mesh accuracy. PA-MPVPE and PA-MPJPE numbers are in $m$.
@@
I2L-MeshNet (Moon and Lee, 2020) 7.4 7.6 0.681 0.973
Pose2Mesh (Choi et al., 2020) 7.7 7.8 0.674 0.969
I2UV-HandNet (Chen et al., 2021a) 6.7 6.9 0.707 0.977
METRO (Lin et al., 2021a) 6.5 6.3 0.731 0.984
Tang et al. (Tang et al., 2021) 6.7 6.7 0.724 0.981
Mesh Graphormer (Lin et al., 2021b) 5.9 6.0 0.764 0.986
MobRecon (Chen et al., 2022) 5.7 5.8 0.784 0.986
AMVUR (Jiang et al., 2023) 6.2 6.1 0.767 0.987
HaMeR (Pavlakos et al., 2024) 6.0 5.7 0.785 0.990
HaMeR (Reproduced) 6.2 5.9 0.774 0.989
@@
Finally, we show that Rodrigues Networks extend beyond robots to animated characters. Specifically, we apply our method to 3D hand reconstruction from single-view RGB images, which involves predicting the rotations and positions of hand joints based on the kinematic structure defined by the MANO (Romero et al., 2022) hand kinematics.
Our network builds upon HaMeR (Pavlakos et al., 2024) by replacing its vanilla transformer with the proposed Rodrigues Network (with modifications to suit MANO’s configuration representation). Additionally, we introduce a cross-attention layer following the self-attention layer to enable interactions between joint and link features and the input image tokens. Our network achieves a notable performance improvement while significantly reducing the number of parameters (39.5M vs. ours: 10.7M). Further architectural details are provided in Section A of the supplementary material.
Table 3 presents the quantitative results. We follow the evaluation protocol and metrics established by HaMeR (Pavlakos et al., 2024), and report performance on the FreiHand (Zimmermann et al., 2019) dataset. For reference, results of our reproduced HaMeR model are also included.
@@ 6 Conclusions and discussions
First, while Rodrigues Networks successfully incorporate articulated kinematics as an inductive bias, they do not yet account for the geometry of individual links. In many settings, this information is available and could improve performance on tasks that require precise contact reasoning. Second, the current Neural Rodrigues Operator is restricted to rotational joints. Extending it to also handle translational joints would broaden its applicability. Third, our robot learning experiments focused on imitation learning. Exploring reinforcement learning (RL) scenarios could further test the network’s generality and effectiveness in closed-loop settings.
@@ 7 Acknowledgements
Leonidas Guibas, Yang You and Congyue Deng acknowledge support from the Toyota Research Institute University 2.0 Program, ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a gift from the Flexiv corporation. Haoran Geng acknowledges support from the Berkeley Fellowship Award and the Qualcomm Innovation Fellowship. Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University. Congyue Deng is also supported in part by the Tayebati Postdoctoral Fellowship at the MIT Schwarzman College of Computing. Pieter Abbeel holds concurrent appointments as a professor at UC Berkeley and as an Amazon Scholar, and this paper is not associated with Amazon.
@@
This supplementary material provides additional details to support the experiments presented in the main paper. Section A describes how we adapt the Neural Rodrigues Operator for animated characters, as used in the 3D hand reconstruction experiment. Section B provides additional details on network architecture. Section C outlines our experimental settings, including data collection, training procedures, evaluation protocols, and all training hyperparameters. Section D details the architecture and configurations of both baseline models and our proposed method across all experiments. These details are provided to facilitate reproducibility of our results. We will release our code, datasets, and pretrained checkpoints upon acceptance to further support the community in reproducing and building upon our work.
Additionally, Section E presents further experiments designed to investigate the properties of our model and provide additional support for the claims made in the main paper. Finally, Section F describes our CUDA implementation of the Multi-Channel Rodrigues Operator, which accelerates training and inference.